# Daisy-TTS: Simulating Wider Spectrum of Emotions via Prosody Embedding Decomposition

### v1

@@ 5.3.1 Primary Emotions Evaluation
Interestingly, if we observe the ground truth’s emotion perceivability, it shows that the ground-truth emotions are not easily distinguishable by humans either. This actually conforms with recent insights in Troiano et al. (2021) on the frequent inconfidence of humans in rating emotional tasks. We will further discuss this in Section 6.1.
@@ 5.3.3 Intensity-level Evaluation
In emotion perception test, we observe that by increasing the intensity-level, the emotion perceivability is improved as well, with Daisy-TTS achieves higher improvement across emotions except for sadness, which conforms to previous tests. This behavior can be explained intuitively from the assumption of emotions would be perceived more confidently if expressed with stronger intensity Troiano et al. (2021).
@@ 6.1 Human Confusion across Primary Emotions
As reported in Section 5.3.1, we found it intriguing that even the emotions in ground truth samples are not easily distinguishable by humans either. But, we could actually explain this from two perspectives. First, as described in Section 2.1, emotions may vary in their degree of similarity to one another, thus some emotions may be perceived similarly and confused with the others. Second, based on the findings in Section 5.3.3 and Troiano et al. (2021), it might also be due to the intensity-level of the evaluated samples.
@@
Table 3: Cosine Similarity between Primary Emotion Embeddings.
Table 4: Confusion matrix of ground truth.
Table 5: Confusion Matrix of Baseline.
@@
True Label
Sadness 0.56 0.21 0.04 0.19
@@
True Label
Sadness 0.53 0.27 0.06 0.13
@@
True Label
Sadness 0.48 0.32 0.08 0.11
@@
Table 3: Cosine Similarity between Primary Emotion Embeddings.
Table 4: Confusion matrix of ground truth.
Table 5: Confusion Matrix of Baseline.
@@
6.2 Ablation of Emotion Classifier
As our method to simulate emotions relies on the emotion separability of the prosody embeddings, it is worth to examine whether training the model with emotion classifier actually affects the learned embeddings.
Figure 5: (Left Column) Prosody embeddings from training Daisy-TTS without emotion classifier; (Right Column) with emotion classifier. Training without emotion classifier discouraged emotion separability.
Figure 5 shows the comparison between the embeddings with and without the emotion classifier. Absence of emotion classifier does make the embedding space not emotionally-separable. However, it turns out that the embeddings would be separated by the speaker’s identity. But, this speaker separability is retained in model with emotion classifier, albeit locally in each emotion cluster.
Figure 6: Explained variance ratio of the prosody embeddings from the ablation. Uniform variance in the absence of an emotion classifier might disallows meaningful decomposition for emotion simulation.
Furthermore, if we examine the explained variance ratio across prosody prototypes (eigenvectors), as shown in Figure 6, it seems like each eigenvector in model not trained with emotion classifier does not describe as meaningful features as the ones trained with emotion classifier. Thus, disallowing emotion simulation through prosody embedding decomposition described in Section 4.
@@ Appendix A Appendix
A.1 Secondary Emotions in Structural Model of Emotions
Through a series of psychological experiments described in Plutchik (1982), the structural model of emotions derived a list of secondary emotions as mixture of primary emotions. Table 7 shows the secondary emotions and their corresponding primary emotion pairs. We incorporate this secondary emotions formulation into our proposed model, Daisy-TTS, and experiments.
@@ A.1 Secondary Emotions in Structural Model of Emotions
A.2 Crowdsourcing Evaluation Setup

### v2

@@ 5.3.1 Primary Emotions Evaluation
Interestingly, if we observe the ground truth’s emotion perceivability, it shows that the ground-truth emotions are not easily distinguishable by humans either. This actually conforms with recent insights in (Troiano et al., 2021) on the frequent inconfidence of humans in rating emotional tasks. We will further discuss this in Section 6.1.
@@ 5.3.3 Intensity-level Evaluation
In emotion perception test, we observe that by increasing the intensity-level, the emotion perceivability is improved as well, with Daisy-TTS achieves higher improvement across emotions except for sadness, which conforms to previous tests. This behavior can be explained intuitively from the assumption of emotions would be perceived more confidently if expressed with stronger intensity (Troiano et al., 2021).
@@ 6.1 Human Confusion across Primary Emotions
As reported in Section 5.3.1, we found it intriguing that even the emotions in ground truth samples are not easily distinguishable by humans either. But, we could actually explain this from two perspectives. First, as described in Section 2.1, emotions may vary in their degree of similarity to one another, thus some emotions may be perceived similarly and confused with the others. Second, based on the findings in Section 5.3.3 and (Troiano et al., 2021), it might also be due to the intensity-level of the evaluated samples.
@@
Table 3: Cosine Similarity between Primary Emotion Embeddings.
@@
True Label Sadness 0.56 0.21 0.04 0.19
@@
True Label Sadness 0.53 0.27 0.06 0.13
@@
True Label Sadness 0.48 0.32 0.08 0.11
@@
6.2 Role of Emotion Discriminator
As our method to simulate emotions relies on the emotion separability of the prosody embeddings, we examine whether training with emotion discriminator actually affects the embeddings.
Figure 5: (Left Column) Prosody embeddings from training Daisy-TTS without emotion discriminator; (Right Column) with emotion discriminator. Training without emotion discriminator discouraged emotion separability.
Figure 5 shows the comparison between the embeddings with and without the emotion discriminator. Absence of emotion discriminator does make the embedding space not emotionally-separable. However, it turns out that the embeddings would be separated by the speaker’s identity. But, this speaker separability is retained in model with emotion discriminator, albeit locally in each emotion cluster.
Figure 6: Explained variance ratio of the prosody embeddings from the ablation. Uniform variance in the absence of an emotion discriminator might disallows meaningful decomposition for emotion simulation.
Furthermore, if we examine the explained variance ratio across prosody prototypes (eigenvectors), as shown in Figure 6, it seems like each eigenvector in model not trained with emotion discriminator does not describe as meaningful features as the ones trained with emotion discriminator. Thus, disallowing emotion simulation through prosody embedding decomposition described in Section 4.
@@ Appendix A Appendix
A.1 Evaluation Results (Cont’d)
Speech Naturalness (MOS; 95% CI) $\uparrow$
Bittersweet Disappointment Outrage
Ground Truth - - -
Baseline 3.433±0.177 3.256±0.221 3.189±0.234
Daisy-TTS 3.667±0.183 3.911±0.168 3.844±0.185
Emotion Perceivability (Acc.) $\uparrow$
Bittersweet Disappointment Outrage
Ground Truth - - -
Baseline 0.211 0.177 0.188
Daisy-TTS 0.288 0.100 0.233
Table 7: Results of MOS and emotion perception tests on primary and secondary emotions, continued from Table 1.
A.2 Secondary Emotions in Structural Model of Emotions
Through a series of psychological experiments described in (Plutchik, 1982), the structural model of emotions derived a list of secondary emotions as mixture of primary emotions. Table 8 shows the secondary emotions and their corresponding primary emotion pairs. We incorporate this secondary emotions formulation into our proposed model, Daisy-TTS, and experiments.
@@ A.2 Secondary Emotions in Structural Model of Emotions
A.3 Crowdsourcing Evaluation Setup